OpenStudy (anonymous):

Find the vertex, focus, directrix, and focal width of the parabola. -1/20x^2=y PLEASE

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

so the parabola has an "x" SQUARED component, so we'd need to use the "focus form" for the parabola of \(\bf (x-h)^2=4p(y-k)\) (h, k) = vertex point p = distance from the vertex to the focus

OpenStudy (jdoe0001):

so let's take a peek at yours \(\bf -\cfrac{1}{2}x^2 = y \implies -\cfrac{1}{2}(x-0)^2 = (y-0)\\ \textit{notice the (h, k) vertex}\\ \textit{notice the "p" distance } 4p = -\cfrac{1}{2}\)

OpenStudy (jdoe0001):

hmm, wait a second... I got ... . a bit off there, lemme rewrite that

OpenStudy (anonymous):

so p =-1/8?

OpenStudy (jdoe0001):

well, from that yes, but the focus form isn't simplified

OpenStudy (jdoe0001):

\(\bf -\cfrac{1}{2}x^2 = y \implies -\cfrac{1}{2}(x-0)^2 = (y-0) \implies (x-0)^2 = -2(y-0)\\ \textit{notice the (h, k) vertex}\\ \textit{notice the "p" distance } 4p = -2\)

OpenStudy (anonymous):

ahh okay so p is -1/2

OpenStudy (jdoe0001):

yes, so the vertex is the origin the focus is negative, means the parabola opens downward, with a hump the focus is 1/2 from there over the y-axis DOWNWARDS, that is (0, -1/2) directrix distance to the vertex is the same as the focus, 1/2, in the opposite direction

OpenStudy (jdoe0001):

and the focus with is 4P

OpenStudy (jdoe0001):

width rather

OpenStudy (anonymous):

so how do i find what the focal width is??

OpenStudy (jdoe0001):

is 4p

OpenStudy (jdoe0001):

just bear in mind that "p" is neither negative nor positive, just a distance unit from the vertex

OpenStudy (jdoe0001):

so the focal width will be 4p => 4( 1/2 )

OpenStudy (anonymous):

I was thinking that but I got focal width as 2 and that wasn't on my options:/

OpenStudy (anonymous):

that's why i'm a bit confused

OpenStudy (jdoe0001):

what options do you have?